# On diffusion-based generative models and their error bounds: The log-concave case with full convergence estimates

### v4

@@ 4.2 Assumptions on the Data Distribution
The vast majority of the results available in the literature are in KL divergence and TV distance. For two general data distributions $\mu$ and $\nu$, there is no known relationship between their KL divergence and their $W_{2}$. However, for strongly log-concave data distributions, a bound on the Wasserstein distance of order two in terms of KL divergence follows from an extension of Talagrand’s inequality (Gozlan and Léonard, 2010, Corollary 7.2).
@@
Some convergence results in different metrics can be deduced under the following types of assumptions. Convergence bounds in Wasserstein distance of order one with exponential complexity has been obtained in (Bortoli, 2022) under the so-called manifold hypothesis, namely assuming that the target distribution is supported on a lower-dimensional manifold or is given by some empirical distribution. Moreover, the results in TV distance in (Lee et al., 2022) and in KL divergence (Yang and Wibisono, 2023) assumed that the data distribution satisfies a logarithmic Sobolev inequality and the score function is Lipschitz resulting in convergence bounds characterized by polynomial complexity. By replacing the requirement that the data distribution satisfies a functional inequality with the assumption that $\pi_{\text{D}}$ has finite KL divergence with respect to the standard Gaussian and by assuming that the score function for the forward process is Lipschitz, the authors in (Chen et al., 2023d) managed to derive bounds in TV distance which scale polynomially in all the problem parameters. By requiring only the Lipschitzness of the score at the initial time instead of the whole trajectory, the authors in (Chen et al., 2023a, Theorem 2.5) managed to show, using an exponentially decreasing then linear step size, convergence bounds in KL divergence with quadratic dimensional dependence and logarithmic complexity in the Lipschitz constant. In the work by (Benton et al., 2024), the authors provide convergence bounds in KL divergence that are linear in the data dimension up to logarithmic factors.
@@
We summarise the results of Chen et al. (2023d, a); Benton et al. (2024) and compare them to ours in Table 2. A careful examination of (Chen et al., 2023a, Proof of Theorem 2.2) and (Benton et al., 2024, Proof of Theorem 1 and Corollary 1) reveals that the authors require to have uniqueness of solutions for the backward SDE (3). Therefore, additional integrability conditions on the score function depending on the type of uniqueness of solutions considered are still needed.
@@
We emphasize that in Theorem 3.10, we do not assume the score function to be Lipschitz continuous with a uniformly bounded Lipschitz constant. This is particularly useful for future work in nonconvex settings, where the upper bound estimates will be independent of the (potentially large) Lipschitz constant of the score function, which could otherwise hide additional dimensional dependencies. The requirement $\int_{\epsilon}^{T}|\nabla\log p_{t}(0)|^{2}\text{d}t<\infty$ in Assumption 2 is weaker than the Lipschiz assumption on the score function, but it is still difficult to verify in practical applications.
@@
As pointed out in (Chen et al., 2023d, Section 4), some type of log-concavity assumption on the data distribution is needed to derive polynomial convergence rates in 2-Wasserstein distance. This justifies the need for our Assumption 2 under which bounds linear in the data dimension are derived in Theorem 3.10. The concurrent result (Gao et al., 2023) has a similar assumption.
Optimization  estimate Regularity  condition Metric Complexity Result
None $\forall t$, $\nabla\log p_{t}$, $L$-Lipschitz $\text{TV}(\mathscr{L}(\widehat{Y}_{K+1}),\pi_{\text{D}})$ $\tilde{O}\left(\frac{ML^{2}}{\delta^{2}}\right)$ (Chen et al., 2023d, Theorem 2)
None $\forall t$, $\nabla\log p_{t}$ $L$-Lipschitz $\sqrt{\text{KL}(\mathscr{L}(\widehat{Y}_{K+1}^{\text{EI}})||\pi_{\text{D}})}$ $\tilde{O}\left(\frac{ML^{2}}{\delta^{2}}\right)$ (Chen et al., 2023a, Theorem 2.1)
None Integrability conditions on $\nabla\log p_{t}$ for the uniqueness of $(Y_{t})_{t\in[0,T]}$ $\sqrt{\text{KL}(\mathscr{L}(\widehat{Y}_{K}^{\text{EI}})||\pi_{\text{D}}^{\underline{\epsilon}})}$ $\tilde{O}\left(\frac{M^{2}\log^{2}(1/\epsilon)}{\delta^{2}}\right)$ (Chen et al., 2023a, Theorem 2.2)
None Integrability conditions on $\nabla\log p_{t}$ for the uniqueness of $(Y_{t})_{t\in[0,T]}$ $\sqrt{\text{KL}(\mathscr{L}(\widehat{Y}_{K}^{\text{EI}})||\pi_{\text{D}}^{\underline{\epsilon}})}$ $\tilde{O}\left(\frac{M\log^{2}(1/\epsilon)}{\delta^{2}}\right)$ (Benton et al., 2024, Corollary 1)
Yes $\pi_{\text{D}}\sim\mathscr{N}(\mu,I_{M})$ $W_{2}(\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})$ $\tilde{O}(\sqrt{M}\log(\delta^{-1}))$ Theorem 3.1
Yes Assumption 2 $W_{2}(\mathscr{L}(\widehat{Y}_{K}^{\text{EM}}),\pi_{\mathsf{D}})$ $\tilde{O}(M^{1/\alpha}(\widetilde{\varepsilon}_{\text{AL}}\epsilon)^{1/(2\alpha)}\log(\delta^{-1}))$ Theorem 3.10
@@
Table 2: Summary of previous bounds and our results. Bounds expressed in terms of the number of steps required to guarantee an error of at most $\delta$ in the stated metric. The variance -$\underline{\epsilon}$ Gaussian perturbation of the target data distribution is denoted by $\pi_{\text{D}}^{\underline{\epsilon}}$. All the bounds assume that $\pi_{\text{D}}$ has finite second moments.
Acknowledgments.
@@
This work has been supported by The Alan Turing Institute through the Theory and Methods Challenge Fortnights event “Accelerating generative models and nonconvex optimisation”, which took place at The Alan Turing Institute headquarters. This work was made possible by a Research-in-Groups programme funded by the International Centre for Mathematical Sciences, Edinburgh. This work is supported by the Guangzhou-HKUST(GZ) Joint Funding Program (No. 2024A03J0630). This work has receved funding from the Ministry of Trade, Industry and Energy (MOTIE) and Korea Institute for Advancement of Technology (KIAT) through the International Cooperative R&D program (No.P0025828). S.S. was supported by the Alan Turing Institute under the EPSRC grant EP/N510129/1.
@@ Appendix A Objective Function via Denoising Score Matching
In this section, we show that the objective function $U$ in (6) can be written into (7) using denoising score matching (Vincent, 2011) and the OU representation
@@
where $p_{t|0}$ is the density of the transition kernel associated with (1) and $p_{t,0}$ is the joint density of $X_{t}$ and $X_{0}$. Using (27) in the objective function given in (6) and the OU representation (26), we have
@@
$\begin{split}U(\theta)&=\int_{\epsilon}^{T}\frac{\kappa(t)}{T-\epsilon}\int_{\mathbb{R}^{M}}\left(|\nabla\log p_{t}(x)|^{2}-2\langle\nabla\log p_{t}(x),s(t,\theta,x)\rangle+|s(t,\theta,x)|^{2}\right)\ p_{t}(x)\ \text{d}x\ \text{d}t\\ &=\int_{\epsilon}^{T}\frac{\kappa(t)}{T-\epsilon}\int_{\mathbb{R}^{M}}\int_{\mathbb{R}^{M}}\Bigg{(}|\nabla\log p_{t}(x)|^{2}-2\left\langle\nabla_{x}\log(p_{t|0}(x|\tilde{x})),\ s(t,\theta,x)\right\rangle\\ &\qquad\qquad\qquad\qquad\qquad+|s(t,\theta,x)|^{2}\Bigg{)}p_{t,0}(x,\tilde{x})\ \text{d}x\ \text{d}\tilde{x}\ \text{d}t\end{split}$
$\begin{split}&=\int_{\epsilon}^{T}\frac{\kappa(t)}{T-\epsilon}\int_{\mathbb{R}^{M}}\int_{\mathbb{R}^{M}}|\nabla_{x}\log p_{t|0}(x|\tilde{x})-s(t,\theta,x)|^{2}p_{t,0}(x,\tilde{x})\ \text{d}x\ \text{d}\tilde{x}\ \text{d}t\\ &\quad+\int_{\epsilon}^{T}\frac{\kappa(t)}{T-\epsilon}\int_{\mathbb{R}^{M}}|\nabla\log p_{t}(x)|^{2}p_{t}(x)\ \text{d}x\ \text{d}t\\ &\quad-\int_{\epsilon}^{T}\frac{\kappa(t)}{T-\epsilon}\int_{\mathbb{R}^{M}}\int_{\mathbb{R}^{M}}|\nabla_{x}\log p_{t|0}(x|\tilde{x})|^{2}p_{t,0}(x,\tilde{x})\ \text{d}x\ \text{d}\tilde{x}\ \text{d}t\\ &=\mathbb{E}[\kappa(\tau)|\sigma_{\tau}^{-1}Z+s(\tau,\theta,m_{\tau}X_{0}+\sigma_{\tau}Z)|^{2}]+\int_{\epsilon}^{T}\frac{\kappa(t)}{T-\epsilon}\int_{\mathbb{R}^{M}}|\nabla\log p_{t}(x)|^{2}p_{t}(x)\ \text{d}x\ \text{d}t\\ &\quad-\int_{\epsilon}^{T}\frac{\kappa(t)}{T-\epsilon}\int_{\mathbb{R}^{M}}\int_{\mathbb{R}^{M}}|\nabla_{x}\log p_{t|0}(x|\tilde{x})|^{2}p_{t,0}(x,\tilde{x})\ \text{d}x\ \text{d}\tilde{x}\ \text{d}t.\end{split}$ (28)
@@
We emphasize that we may choose to restrict $t\in[\epsilon,T]$ with $\epsilon\in(0,1)$ to prevent the divergence of the integrals on the right-hand side of (28).

### v5

@@ 4.2 Assumptions on the Data Distribution
The vast majority of the results available in the literature are in KL divergence and TV distance. For two general data distributions $\mu$ and $\nu$, there is no known relationship between their KL divergence and their $W_{2}$. However, for strongly log-concave data distributions, a bound on the Wasserstein distance of order two in terms of KL divergence follows from an extension of Talagrand’s inequality Gozlan & Léonard (2010, Corollary 7.2).
@@
Some convergence results in different metrics can be deduced under the following types of assumptions. Convergence bounds in Wasserstein distance of order one with exponential complexity has been obtained in Bortoli (2022) under the so-called manifold hypothesis, namely assuming that the target distribution is supported on a lower-dimensional manifold or is given by some empirical distribution. Moreover, the results in TV distance in Lee et al. (2022) and in KL divergence Yang & Wibisono (2023) assumed that the data distribution satisfies a logarithmic Sobolev inequality and the score function is Lipschitz resulting in convergence bounds characterized by polynomial complexity. By replacing the requirement that the data distribution satisfies a functional inequality with the assumption that $\pi_{\text{D}}$ has finite KL divergence with respect to the standard Gaussian and by assuming that the score function for the forward process is Lipschitz, the authors in Chen et al. (2023d) managed to derive bounds in TV distance which scale polynomially in all the problem parameters. By requiring only the Lipschitzness of the score at the initial time instead of the whole trajectory, the authors in Chen et al. (2023a, Theorem 2.5) managed to show, using an exponentially decreasing then linear step size, convergence bounds in KL divergence with quadratic dimensional dependence and logarithmic complexity in the Lipschitz constant. In the work by Benton et al. (2024), the authors provide convergence bounds in KL divergence that are linear in the data dimension, up to logarithmic factors, using early stopping and assuming finite second moments of the data distribution. A careful examination of Benton et al. (2024, Proof of Theorem 1 and Corollary 1) reveals that the authors still require the uniqueness of solutions for the backward SDE 3. For instance, either measurability and boundedness of the drift (Stroock & Varadhan, 1997) or integrability of the drift (Röckner & Zhao, 2023) should be imposed if the uniqueness of weak solutions is considered. Therefore, additional conditions on the score function, depending on the type of uniqueness of solutions considered, are still needed in their bounds in Theorem 1 and Corollary 1.
Assuming the finiteness of the second moment of the data distribution and using an EI discretization scheme with constant and exponentially decreasing step sizes, the authors in Conforti et al. (2025, Corollary 2.4) derive a KL divergence bound with early stopping, which scales linearly in the data dimension up to logarithmic factors. In terms of dependence of the data dimension, this bound is slightly worse than the bound in $W_{2}$ provided in Remark 12, which does not include the logarithmic dependence of $M$. Bounds in KL without early stopping have been derived in Conforti et al. (2025) for data distributions with finite Fisher information with respect to the standard Gaussian distribution. The bounds in Conforti et al. (2025, Theorem 2.1 and 2.2) scale linearly in the Fisher information when an EI discretization scheme with constant step size is used and logarithmically in the Fisher information when an exponential-then-constant step size Conforti et al. (2025, Theorem 2.3) is chosen. We note that Conforti et al. (2025, Theorem 2.1 and 2.2) cannot achieve the optimal rate of convergence of Theorem 1 in the motivating example.
@@
We summarise the results of Chen et al. (2023a); Benton et al. (2024); Conforti et al. (2025) and compare them to ours in Table 2 and Table 3, making the distinction based on whether the early stopping rule is applied. In addition, a careful examination of Chen et al. (2023a, Proof of Theorem 2.2) reveals that the authors require the uniqueness of solutions for the backward SDE 3. For instance, when strong solutions are considered, this implies that the score function should be (at least) monotone in the space variable, and a suitable integrability condition in $t$ (similar to our Assumption 2) is still needed to guarantee uniqueness of the solution (see, e.g., Krylov (1991, Theorem 1)). For weak solutions, we refer to the discussion of Benton et al. (2024, Proof of Theorem 1 and Corollary 1) above.
@@
We emphasize that in Theorem 10, we do not assume the score function to be Lipschitz continuous with a uniformly bounded Lipschitz constant. This is particularly useful for future work in nonconvex settings, where the upper bound estimates will be independent of the (potentially large) Lipschitz constant of the score function, which could otherwise hide additional dimensional dependencies. The requirement $\int_{\epsilon}^{T}|\nabla\log p_{t}(0)|^{2}\text{d}t<\infty$ in Assumption 2 is weaker than the Lipschitz assumption on the score function, but it is still difficult to verify in practical applications.
@@
As pointed out in Chen et al. (2023d, Section 4), some type of log-concavity assumption on the data distribution is needed to derive polynomial convergence rates in 2-Wasserstein distance. This justifies the need for our Assumption 2. The concurrent result Gao et al. (2023) has a similar assumption. For completeness, we also mention that the results in the same metric, which have appeared after the first version of our preprint on arXiv (e.g., Tang & Zhao (2024); Strasman et al. (2025)), continue to assume strong log-concavity of the data distribution.
@@
Table 2: Summary of previous bounds without early stopping and our result in Theorem 1. Bounds expressed in terms of the number of steps required to guarantee an error of at most $\delta$ in the stated metric. The relative Fisher information of the target $\pi_{\text{D}}$ against standard Gaussian measure $\pi_{\infty}$ is denoted by $\text{FI}(\pi_{\text{D}}|\pi_{\infty})$. All the bounds assume that $\pi_{\text{D}}$ has finite second moments.
Optimization  problem solved Regularity  condition Metric Complexity Reference
No $\forall t$, $\nabla\log p_{t}$ $L$-Lipschitz $\sqrt{\text{KL}(\mathcal{L}(\widehat{Y}_{K+1}^{\text{EI}})||\pi_{\text{D}})}$ $\tilde{O}\left(\frac{ML^{2}}{\delta}\right)$ Chen et al. (2023a, Theorem 2.1)
No Conforti et al. (2025, H2), i.e. $\text{FI}(\pi_{\text{D}}|\pi_{\infty})<\infty$ $\sqrt{\text{KL}(\mathcal{L}(\widehat{Y}_{K+1}^{\text{EI}})||\pi_{\text{D}})}$ $\tilde{O}\left(\sqrt{\frac{M+\mathbb{E}[|X_{0}|^{2}]}{\delta}\log^{2}(\mathfrak{L})}\right)$, with $\mathfrak{L}:=M^{-1}\text{FI}(\pi_{\text{D}}|\pi_{\infty})$ Conforti et al. (2025, Theorem 2.3)
Yes $\pi_{\text{D}}\sim\mathcal{N}(\mu,I_{M})$ $W_{2}(\mathcal{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})$ $\tilde{O}(\frac{\sqrt{M}}{\delta})$ Theorem 1
Table 3: Summary of previous bounds with early stopping and our results in Remark 12. The results in Chen et al. (2023a, Theorem 2.2), Benton et al. (2024, Theorem 1), and Conforti et al. (2025, Corollary 2.4) are stated for the smoothed version of the data, denoted by $\pi^{\epsilon}_{\text{D}}$ and using the score approximation assumption 26 with $\underline{\varepsilon}$. Therefore, an additional error should be added to their bounds, as the distance between $\pi^{\epsilon}_{\text{D}}$ and $\pi_{\text{D}}$ scales with $\sqrt{M}$ in $W_{2}$ (see Appendix D.2).
Assumption on the data Metric Error bound Reference
Finite second moments of $\pi_{\text{D}}$ and 26. See Section 4.2 for conditions on $\nabla\log p_{t}$ for the uniqueness of $Y_{t}$. $\sqrt{\text{KL}(\mathcal{L}(\widehat{Y}_{K}^{\text{EI}})||\pi^{\epsilon}_{\text{D}})}$ $\sqrt{(\mathbb{E}[|X_{0}|^{2}]+M)e^{-T}}+\sqrt{T\underline{\varepsilon}}+M(T+\log\epsilon^{-1})/\sqrt{N}$, (+ additional bounds between $\pi^{\epsilon}_{\text{D}}$ and $\pi_{\text{D}}$) Chen et al. (2023a, Theorem 2.2)
Finite second moments of $\pi_{\text{D}}$ and 26. See Section 4.2 for conditions on $\nabla\log p_{t}$ for the uniqueness of $Y_{t}$. $\sqrt{\text{KL}(\mathcal{L}(\widehat{Y}_{K}^{\text{EI}})||\pi^{\epsilon}_{\text{D}})}$ $\sqrt{\underline{\varepsilon}}+\sqrt{\tilde{O}(M/N)}+\sqrt{Me^{-2T}}$, (+ additional bounds between $\pi^{\epsilon}_{\text{D}}$ and $\pi_{\text{D}}$) Benton et al. (2024, Theorem 1)
Finite second moments of $\pi_{\text{D}}$ and 26. $\sqrt{\text{KL}(\mathcal{L}(\widehat{Y}_{K}^{\text{EI}})||\pi^{\epsilon}_{\text{D}})}$ $\sqrt{(\mathbb{E}[|X_{0}|^{2}]+M)e^{-T}}+\sqrt{T\underline{\varepsilon}}+[c[(M+\mathbb{E}[|X_{0}|^{2}])(\log(M+\mathbb{E}[|X_{0}|^{2}])+\log(\epsilon^{-1})+1)]]^{1/2}$, with $c,\epsilon\in(0,1/2]$ (+ additional bounds between $\pi^{\epsilon}_{\text{D}}$ and $\pi_{\text{D}}$) Conforti et al. (2025, Corollary 2.4)
Assumption 2 and Assumption 4. $W_{2}(\mathcal{L}(\widehat{Y}_{K}^{\text{EM}}),\pi_{\mathsf{D}})$ $O(\sqrt{M})\sqrt{\epsilon}+O(\sqrt{M})e^{-2\widehat{L}_{\text{MO}}(T-\epsilon)-\epsilon}+O(e^{(1+\zeta-2\widehat{L}_{\text{MO}})(T-\epsilon)})\sqrt{\varepsilon_{\text{SN}}}+O(\sqrt{M}e^{T^{2\alpha+1}}T^{2\alpha+1}\widetilde{\varepsilon}^{1/2}_{\text{AL}})\gamma^{1/2}$ Remark 12
Acknowledgments
@@
This work has been supported by The Alan Turing Institute through the Theory and Methods Challenge Fortnights event “Accelerating generative models and nonconvex optimisation”, which took place at The Alan Turing Institute headquarters. This work was made possible by a Research-in-Groups programme funded by the International Centre for Mathematical Sciences, Edinburgh. This work is supported by the Guangzhou-HKUST(GZ) Joint Funding Programs (No. 2024A03J0630 and No. 2025A03J3322). This work has received funding from the Ministry of Trade, Industry and Energy (MOTIE) and Korea Institute for Advancement of Technology (KIAT) through the International Cooperative R&D program (No.P0025828). S.S. was supported by the Alan Turing Institute under the EPSRC grant EP/N510129/1.
Appendix
@@ Appendix A Objective Function via Denoising Score Matching
In this section, we show that the objective function $U$ in 6 can be written into 7 using denoising score matching (Vincent, 2011) and the OU representation
@@
where $p_{t|0}$ is the density of the transition kernel associated with 1 and $p_{t,0}$ is the joint density of $X_{t}$ and $X_{0}$. Using 31 in the objective function given in 6 and the OU representation 30, we have
@@
$\begin{split}U(\theta)&=\int_{\epsilon}^{T}\frac{\kappa(t)}{T-\epsilon}\int_{\mathbb{R}^{M}}\left(|\nabla\log p_{t}(x)|^{2}-2\langle\nabla\log p_{t}(x),s(t,\theta,x)\rangle+|s(t,\theta,x)|^{2}\right)\ p_{t}(x)\ \text{d}x\ \text{d}t\\ &=\int_{\epsilon}^{T}\frac{\kappa(t)}{T-\epsilon}\int_{\mathbb{R}^{M}}\int_{\mathbb{R}^{M}}\Bigg{(}|\nabla\log p_{t}(x)|^{2}-2\left\langle\nabla_{x}\log(p_{t|0}(x|\tilde{x})),\ s(t,\theta,x)\right\rangle\\ &\qquad\qquad\qquad\qquad\qquad+|s(t,\theta,x)|^{2}\Bigg{)}p_{t,0}(x,\tilde{x})\ \text{d}x\ \text{d}\tilde{x}\ \text{d}t\\ &=\int_{\epsilon}^{T}\frac{\kappa(t)}{T-\epsilon}\int_{\mathbb{R}^{M}}\int_{\mathbb{R}^{M}}|\nabla_{x}\log p_{t|0}(x|\tilde{x})-s(t,\theta,x)|^{2}p_{t,0}(x,\tilde{x})\ \text{d}x\ \text{d}\tilde{x}\ \text{d}t\\ &\quad+\int_{\epsilon}^{T}\frac{\kappa(t)}{T-\epsilon}\int_{\mathbb{R}^{M}}|\nabla\log p_{t}(x)|^{2}p_{t}(x)\ \text{d}x\ \text{d}t\\ &\quad-\int_{\epsilon}^{T}\frac{\kappa(t)}{T-\epsilon}\int_{\mathbb{R}^{M}}\int_{\mathbb{R}^{M}}|\nabla_{x}\log p_{t|0}(x|\tilde{x})|^{2}p_{t,0}(x,\tilde{x})\ \text{d}x\ \text{d}\tilde{x}\ \text{d}t\\ &=\mathbb{E}[\kappa(\tau)|\sigma_{\tau}^{-1}Z+s(\tau,\theta,m_{\tau}X_{0}+\sigma_{\tau}Z)|^{2}]+\int_{\epsilon}^{T}\frac{\kappa(t)}{T-\epsilon}\int_{\mathbb{R}^{M}}|\nabla\log p_{t}(x)|^{2}p_{t}(x)\ \text{d}x\ \text{d}t\\ &\quad-\int_{\epsilon}^{T}\frac{\kappa(t)}{T-\epsilon}\int_{\mathbb{R}^{M}}\int_{\mathbb{R}^{M}}|\nabla_{x}\log p_{t|0}(x|\tilde{x})|^{2}p_{t,0}(x,\tilde{x})\ \text{d}x\ \text{d}\tilde{x}\ \text{d}t.\end{split}$ (32)
@@
We emphasize that we may choose to restrict $t\in[\epsilon,T]$ with $\epsilon\in(0,1)$ to prevent the divergence of the integrals on the right-hand side of 32.
Appendix B Additional Discussions about Assumption 4
In this section, we provide additional details about Remark 8, in which we used the multivariate logistic distribution (Malik & Abraham, 1973)
$\begin{split}g(x)&=M!\ \exp\left(-\sum_{k=1}^{M}x_{k}\right)\left(1+\sum_{k=1}^{M}\exp(-x_{k})\right)^{-(M+1)},\qquad k=1,\dots,M,\quad-\infty<x_{k}<\infty,\end{split}$
and derived its score function in 23. The Hessian of $\log(g(x))$ is
$\begin{split}\frac{\partial^{2}}{\partial x_{k}^{2}}\log(g(x))&=-(M+1)\left(\frac{\exp(-x_{k})}{(1+\sum_{k=1}^{M}\exp(-x_{k}))}-\frac{\exp(-2x_{k})}{(1+\sum_{k=1}^{M}\exp(-x_{k}))^{2}}\right),\\ \frac{\partial^{2}}{\partial x_{k}\partial x_{j}}\log(g(x))&=-(M+1)\left(\frac{\exp(-x_{k})\exp(-x_{j})}{(1+\sum_{k=1}^{M}\exp(-x_{k}))^{2}}\right),\qquad j,k=1,\dots,M,\quad j\neq k,\end{split}$ (33)
Since the Hessian 33 is bounded, we can conclude that the score function of the multivariate logistic distribution defined in 23 is Lipschitz continuous. By using the same arguments as in Remark 8, we can conclude that $\nabla\log p_{t}(x)$ defined in 22 is Lipschitz.
Using Assumption 3.b and Assumption 1, we have
$\begin{split}&\mathbb{E}\int_{0}^{T-\epsilon}|\nabla\log p_{T-r}(Y_{r}^{\text{aux}})-s(T-r,\hat{\theta},Y_{r}^{\text{aux}})|^{2}\ \text{d}r\\ &\leq 2\mathbb{E}\int_{0}^{T-\epsilon}|\nabla\log p_{T-r}(Y_{r}^{\text{aux}})-s(T-r,\theta^{*},Y_{r}^{\text{aux}})|^{2}\ \mathrm{d}r+2\mathsf{K}_{2}^{2}\int_{0}^{T-\epsilon}(1+2|T-r|^{\alpha})^{2}\ \mathbb{E}[|\hat{\theta}-\theta^{*}|^{2}]\ \mathrm{d}r\\ &\leq 2\mathbb{E}\int_{0}^{T-\epsilon}|\nabla\log p_{T-r}(Y_{r}^{\text{aux}})-s(T-r,\theta^{*},Y_{r}^{\text{aux}})|^{2}\ \mathrm{d}r+4\mathsf{K}_{2}^{2}(T-\epsilon)(1+4T^{2\alpha})\widetilde{\varepsilon}_{\text{AL}}<\varepsilon_{\text{SN}},\end{split}$
where the error of the first term on the right-hand side above is expected to be small since the Lipschitz $\nabla\log p_{t}(x)$ is approximated by the function $s$ satisfying Assumption 1. This satisfies Assumption 4.
In addition, we remark that the motivating example in section 3.1 is a special case of the general setting. Indeed, Assumption 1 is satisfied due to Lemma 14 with $\hat{\theta}=\theta_{n}^{\lambda}$ being the $n$th-iterate of the SGLD algorithm 14 and $\theta^{*}:=\mu$. Assumption 2 is satisfied by the score function 12. Assumption 3.b is satisfied by the approximating function 13 with $d=M$, $\alpha=1$, $D_{1}(\theta,\bar{\theta}):=|\bar{\theta}|$, $D_{2}(t,\bar{t}):=e^{-t}$, $D_{3}(t,\bar{t}):=1$, for any $\theta$, $\bar{\theta}\in\mathbb{R}^{d}$, $t$, $\bar{t}\in[0,T]$ and $\mathsf{K}_{1}=\mathsf{K}_{2}=\mathsf{K}_{3}=\mathsf{K}_{4}=1$. Furthermore, both functions satisfy Assumption 4 with $d=M$. Indeed, by Assumption 1 with $\theta^{*}=\mu$, we have